# An Interface Method for Co-simulation of EMT Model and Shifted Frequency EMT Model Based on Rotational Invariance Techniques

### v1

@@ Abstract
The shifted frequency-based electromagnetic transient (SFEMT) simulation has greatly improved the computational efficiency of traditional electromagnetic transient (EMT) simulation for the ac grid. This letter proposes a novel interface for the co-simulation of the SFEMT model and the traditional EMT model. The general form of SFEMT modeling and the principle of analytical signal construction are first derived. Then, an interface for the co-simulation of EMT and SFEMT simulation is proposed based on rotational invariance techniques. Theoretical analyses and test results demonstrate the effectiveness of the proposed method.
@@ I Introduction
For modern power system, the electromagnetic transient (EMT) is indispensable. In the EMT simulation, a very small time step is adopted for the whole system to capture the dynamics with the minimum time-scale, resulting in low computational efficiency. In fact, the dynamics of a large ac/dc power grid contains transients of time-scales from microseconds to milliseconds. Therefore, it is meaningful to develop a multi-scale simulation method and adopt different modeling methods and time steps for different subsystems.
An idea of multi-rate EMT simulation with large step sizes for the ac grids and small step sizes for dc grids has been proposed for more than twenty years which improves the power system EMT simulation efficiency. However, for traditional EMT simulation of ac grids, the time step-size is still small. To improve the efficiency, the shifted frequency-based EMT (SFEMT) simulation with a much larger step size is proposed [1, 2]. Then, a co-simulation based on SFMET simulation and traditional EMT simulation is proposed in [3, 4] and the interface is designed. In the co-simulation, the ac grid with slow dynamics is simulated by SFEMT simulation and the dc grid with fast dynamics is calculated by traditional EMT simulation. However, the existing interface methods for co-simulation of SFEMT simulation and traditional EMT simulation have a certain degree of accuracy loss. In these interface methods, the complex signals for SFEMT simulation that are generated from real signals in the EMT simulation are not analytical signals that orthogonal to the original real signals.
Aiming at the problem of the existing interfaces, this paper proposes a novel interface between SFEMT and EMT models. First, the general form of SFEMT modeling is derived, laying out the foundation for SFEMT simulation and giving guidance for the co-simulation interface design. Then, an analytical signal construction method based on the estimation of signal parameters via rotational invariance techniques (ESPRIT) is proposed. It can accurately identify the frequency, amplitude and phase of each component in an instantaneous signal. According to the frequency spectrum of the instantaneous signal of the EMT simulation, the analytical signal that the SFEMT simulation needs can be constructed. The proposed interface is more accurate than the existing interfaces.
This letter is organized as follows. Section II derives the principle of SFEMT modeling and analytic signal construction. The interface based on ESPRIT is designed and studied in Sections III and IV. Section V concludes.
II Exploration of the Principle of SFEMT Modeling
@@ II-A General Form of SFEMT Modeling
where $x(t)$ is the state variable, $A$ is the coefficient of $x(t)$, and $u(t)$ is the input. In (1), the nonlinear part of the component dynamic is combined into the input $u(t)$, which is generally calculated by means of delay or prediction. For SFEMT modeling, a mathematical transformation $T[\cdot]$ (e.g., Hilbert transform) with the differential property ($T[\frac{\mathrm{d}x(t)}{\mathrm{d}t}]=\frac{\mathrm{d}T[x(t)]}{\mathrm{d}t}$) and linear property ($T[kx(t)]=kT[x(t)]$) is performed on both sides of (1) simultaneously to construct the adjoint system of (1):
@@
With (1)$+\mathrm{j}$(2), the system dynamic equation can be represented by analytical signal:
@@
where $\underline{x}(t)$ and $\underline{u}(t)$ are the analytical signals corresponding to $x(t)$ and $u(t)$, respectively. Then, the frequency-shifting transformation is performed on both sides of (3):
@@
where $X(t)$and $U(t)$ are the complex envelopes corresponding to $x(t)$and $u(t)$, respectively. By discretizing (4), an SFEMT model for the component can be obtained.
II-B Implementation of the Analytical Signal Consctruction
In addition to the two conditions of transformation $T[\cdot]$ in Section II-A, the analytical signal $\underline{u}(t)$ should not contain frequency components on the negative half axis of the spectrum [2] for the use a large time step-size. Otherwise, this negative frequency signal component will become a signal component with a higher frequency than the original signal component after frequency shifting. Thus, it is also necessary to ensure that the input signal $U(t)$ is a low-frequency envelope signal. Therefore, the $T[\cdot]$ used to construct $\underline{u}(t)$ must also satisfy:
@@ II-B Implementation of the Analytical Signal Consctruction
where $\rm sgn(\cdot)$ is the sign function, and $\mathcal{F}(u(t))$is the spectrum of $u(t)$.
According to the above analyses, the construction method of an analytical signal can be designed. First, for an input $u(t)$ with a known analytical expression, the Hilbert transform can be used [1, 2]. However, the convolution of the signal $u(t)$ and $h(t)$ in Hilbert transformation involves calculations from $\infty$- ∞ to $\infty$, which cannot be implemented for signals with unknown analytical expressions. Second, for signals with unknown analytical expressions, the spectrum analysis-based methods (e.g., FFT and ESPRIT) can be considered, which satisfy the (5) and the two conditions in Section II-A.
@@ III Interface Between EMT Model and SFEMT Model Based on ESPRIT
In the muti-area co-simulation, the distributed transmission line model is widely used for the interface between different areas, which can also be used for co-simulation of the EMT and SFEMT models. The difference is that an interface between the real signal and the complex analytical signal is needed which is detailedly elaborated as follows.
@@ III-A Conversion Between Analytical Signal and Real Signal
The real signal can be obtained from the analytical signal by:
@@
where $x(t)$ is the real signal that the EMT model needs and $X(t)$ is the complex envelope signal in the SFEMT model.
The complex envelope signal that the SFEMT model needs can be calculated by:
$X(t)=(x(t)+\mathrm{j}T[x(t)])\mathrm{e}^{-\mathrm{j}\omega_{\rm s}t}$ (7)
Due to the inability to obtain the analytical expression of the voltage and current in EMT simulation, this paper proposes an analytic signal construction method based on the ESPRIT[5], which can accurately extract the spectrum of the signal.
@@ III-B1 Determination of the Number of Signal Components
Assuming the sampling step size of the signal $x(t)$ is $\Delta t$, $x(t)$ can be discretized to $x(j\Delta t)=x_{j}$. For a window with a length of $T_{\rm w}$, which contains $N=2n+1$ samples. The Hankel matrix $\boldsymbol{X}$ composed of signals within this data window can be represented as:
@@
where every two consecutive singular values correspond to a signal component. The singular values corresponding to signal components are much greater than those corresponding to noises. According to this, the number of signal components $m$ can be determined [6].
III-B2 Calculation of the Frequency of Each Signal Component
According to [7], the Hankel matrix $\boldsymbol{X}$ composed of the signal can be decomposed into:
@@ III-B2 Calculation of the Frequency of Each Signal Component
where $z_{i}=\mathrm{e}^{\mathrm{j}2\uppi\frac{f_{i}}{f_{\rm sp}}}$, $z_{i}^{h}=\mathrm{e}^{\mathrm{j}2\uppi h\frac{f_{i}}{f_{\rm sp}}}$, $f_{\rm sp}$ is the sampling frequency, $\boldsymbol{P}$ is a matrix composed of the frequency spectrum of each signal component (referred to as phasor). It is worth noting that $z_{i}$ can be solved based on the matrix pencil techniques [7]. Then, the frequency of each signal is calculated:
@@ III-B3 Calculation of Amplitude and Phase of Each Signal Component
After obtaining the frequency of each signal component, $\boldsymbol{Z}_{\mathrm{R}}$ and $\boldsymbol{Z}_{\rm L}$ in (10) can be calculated. Next, by using the least squares method to solve (10), the phasor matrix $\boldsymbol{P}$ can be obtained:
@@
where $\boldsymbol{Z}_{\mathrm{L}}^{\mathrm{H}}$ is the conjugate transposition of $\boldsymbol{Z}_{\mathrm{L}}$. Then, the amplitude and phase of each signal component can be obtained based on $\boldsymbol{P}$:
${a}_{i}=2\left|{p}_{ii}\right|$ (13)
${\phi}_{i}=\angle{p}_{ii}$
@@ III-B4 Analytical Signal Consctruction
After the frequency, amplitude, and phase of each signal component are all obtained, the imaginary part signal $T\left(x(t)\right)$ (denoted as $\hat{x}(t)$) in $\underline{x}(t)$ can be accurately calculated:
@@
Based on the original real signal and imaginary part signal, the analytical signal can be constructed:
$\underline{x}(t)=x(t)+\mathrm{j}\hat{x}(t)$ (15)
Then, by using frequency shift transformation, the complex envelope signal is obtained:
$X(t)=\underline{x}(t)\mathrm{e}^{-\mathrm{j}2\omega_{\rm s}t}$ (16)
The analytical signal $\underline{x}(t)$ constructed by ESPRIT is a true analytical signal, which only contains positive-frequency components. The frequencies of the complex envelope signal $X(t)$ obtained by frequency shift transformation are much lower than the original signal $x(t)$. On the contrary, the analytical signals constructed by the methods of delay transformation [3] and companion simulation [4] are virtual analytical signals. When there are harmonic components in the system, if the delay transformation [3] or companion simulation methods are used, the complex envelope signal $X(t)$ will contain signal components with higher frequencies than the original signal.
@@ IV Case Studies
[…]
@@
First, the test system is simulated by the co-simulation with the proposed interface, where $\rm S_{1}$ is simulated by EMT simulation and $\rm S_{2}$ is simulated by SFEMT simulation. The time step-sizes for the two subsystems are 20 $\upmu$s and 500 $\upmu$s, respectively. Then, the test system is calculated by using the co-simulation of EMT and SFEMT simulations with the interface methods in [3, 4]. The current of line 26-27 obtained by different methods is shown in Fig. 2.
Figure 2: Phase A current at the sending terminal of lines 26-27.
Due to the integration of wind farm, a subsynchronous oscillation happens and there are interharmonic components in the current besides the fundamental component. It can be found that the co-simulation based on the proposed interface is more accurate than the other two interfaces when the voltage and current signals at the interface contain inter-harmonic components. The reason is that the imaginary part of the analytical signal obtained by delay transformation and companion simulation does not satisfy the orthogonal relationship (5) with the original signal. The analytical signals obtained with these two methods still contain negative frequency components. After frequency shift transformation, the analytical envelope will contain signals with higher frequencies. On the contrary, the analytical signal constructed based on the proposed method only contains positive frequency components. After frequency shift transformation, the frequency of the signal decreases, and the simulation accuracy will be improved.
Figure 3: The amplitude spectrum of analytical signals obtained by different interface methods.
To further illustrate the advantages of the proposed method, the spectra of the analytical signals obtained from the three different interface methods are compared. For the receiving-end current of transmission line 26-28 that was obtained by different methods, the amplitude spectra of its analytical signal are shown in Fig. 3. It can be observed that the spectra of the analytical signals obtained based on delay transformation or companion simulation interfaces contain negative frequency components. However, the spectrum of the analytical signal obtained by the proposed method only contains positive frequency components. After frequency shifting the signal, the envelope signal in the system will become a low-frequency signal, and a larger simulation step size can be used.
@@ V Conclusion
This letter proposes an interface method for the EMT and the SFMET models of power systems. An analytical signal construction method based on ESPRIT is proposed for the interface, which improves the accuracy of co-simulation. Inspired by this letter, many other methods in the signal process area may be introduced into the area of power system co-simulation.

### v2

# An Interface Method for Co-simulation of EMT Model and Shifted Frequency EMT Model Based on Estimation of Signal Parameters via Rotational Invariance Techniques

@@ Abstract
The shifted frequency-based electromagnetic transient (SFEMT) simulation is much more efficient than traditional electromagnetic transient (EMT) simulation for ac grid. This letter proposes a novel interface for the co-simulation of the SFEMT model and the conventional EMT model. The fundamentals of SFEMT modeling are first derived. Then, an interface for the co-simulation of EMT and SFEMT models is proposed based on the estimation of signal parameters via rotational invariance techniques. Theoretical analyses and test results demonstrate the effectiveness of the proposed method.
@@ I Introduction
In traditional electromagnetic transient (EMT) simulation, a small time step is adopted to capture the dynamics with the minimum time-scale in the system, resulting in low computational efficiency. In fact, the dynamics of a large power system model involve multi-scale transients. Therefore, it is meaningful to develop a multi-scale simulation method that adopts different modeling methods and time steps for different subsystems. An idea of multi-rate EMT simulation with large step sizes for the ac grid and small step sizes for power electronics was proposed twenty years ago. However, for the simulation of ac grids, the time step-size is still small.
To improve simulation efficiency, the shifted frequency-based EMT (SFEMT) simulation with a much larger step size is proposed [1, 2]. Then, the co-simulation based on SFMET simulation and traditional EMT simulation is proposed in [3, 4], in which some subsystems of the power system are calculated by SFEMT simulation and the others are calculated by traditional EMT simulation. However, the interfaces for these co-simulation methods have a certain degree of accuracy loss. In these interfaces, the analytical signals for SFEMT simulation that are generated from real signals in the EMT simulation are not true analytical signals that are orthogonal to the original real signals.
Aiming at the problem of the existing interfaces, this paper proposes a novel interface for the SFEMT and EMT models. First, the general form of SFEMT modeling is derived, giving a guideline for the design of co-simulation interface. Then, an interface based on the estimation of signal parameters via rotational invariance techniques (ESPRIT) is proposed. The ESPRIT can accurately calculate the frequency, amplitude and phase of each component in an instantaneous signal with a short data window. According to the frequency, amplitude and phase of the original signal, the analytical signal that the SFEMT simulation needs can be constructed. The proposed interface is more accurate than the existing interfaces.
The letter is organized as follows. Section II derives SFEMT modeling fundamentals. The interface of co-simulation is designed and studied in Sections III and IV. Section V concludes.
II SFEMT Modeling Fundamentals
@@ II-A General Form of SFEMT Modeling
where $x(t)$ is the state variable, $A$ is the coefficient of $x(t)$, and $u(t)$ is the input. In (1), the nonlinear part of the component dynamic is combined into $u(t)$, which is generally calculated by means of delay or prediction. For SFEMT modeling, a mathematical transformation $T[\cdot]$ (e.g., Hilbert transform) with the differential property ($T[\frac{\mathrm{d}x(t)}{\mathrm{d}t}]=\frac{\mathrm{d}T[x(t)]}{\mathrm{d}t}$) and linear property ($T[kx(t)]=kT[x(t)]$) is performed on both sides of (1) simultaneously to construct the adjoint system of (1):
@@
Through (1)$+\mathrm{j}$(2), the system dynamic equation can be represented by analytical signal:
@@
where $\underline{x}(t)$ and $\underline{u}(t)$ are the analytical signals corresponding to $x(t)$ and $u(t)$, respectively. Then, by performing the frequency-shifting transformation on both sides of (3), (4) is derived.
@@
where $X(t)$ and $U(t)$ are the analytical envelopes corresponding to $x(t)$ and $u(t)$, respectively. By discretizing (4), an SFEMT model for the component can be obtained.
II-B Impletation of Analytical Signal Consctruction
To obtain a true analytical signal, the analytical signal $\underline{u}(t)$ should not contain frequency components on the negative half axis of the spectrum [2] for the use of a large time step-size. Otherwise, the signal component with negative frequency will become a signal component with a higher frequency after frequency shifting (e.g., $-65$ Hz to $-115$ Hz). Thus, to ensure that the signal $U(t)$ is a low-frequency envelope signal, the $T[\cdot]$ that is used to construct $\underline{u}(t)$ must also satisfy:
@@ II-B Impletation of Analytical Signal Consctruction
where $\rm sgn(\cdot)$ is the sign function and $\mathcal{F}(u(t))$ is the spectrum of $u(t)$.
According to the above analyses, the guidelines of the analytical signal construction can be proposed. First, for an input $u(t)$ with a known analytical expression, the Hilbert transform can be used [1, 2]. However, the convolution of the signal $u(t)$ and $h(t)$ in Hilbert transformation involves calculations from $\infty$- ∞ to $\infty$, which cannot be implemented for signals with unknown analytical expressions. Then, for signals with unknown analytical expressions, the spectrum analysis-based methods (e.g., ESPRIT) can be considered, which satisfy (5) and the two conditions in Section II-A.
@@ III Interface Between EMT Model and SFEMT Model Based on ESPRIT
In the muti-area co-simulation, the distributed transmission line model is widely used for the interface between different areas, which can also be used for co-simulation of the EMT and SFEMT models. The difference is that an interface between the real signal and the analytical envelope signal is needed in the co-simulation of the EMT and SFEMT models, which is detailedly elaborated as follows.
@@ III-A Conversion Between Analytical Signal and Real Signal
The real signal can be obtained from the analytical envelope signal by:
@@
where $x(t)$ is the real signal that the EMT model needs and $X(t)$ is the analytical envelope signal in the SFEMT model.
The analytical envelope signal that the SFEMT model needs can be calculated by:
$X(t)=\underline{x}(t)\mathrm{e}^{-\mathrm{j}\omega_{\rm s}t}=(x(t)+\mathrm{j}T[x(t)])\mathrm{e}^{-\mathrm{j}\omega_{\rm s}t}$ (7)
Due to the inability to obtain the analytical expression of the voltage and current in EMT simulation, this paper proposes an analytic signal construction method based on the ESPRIT, which can accurately extract the spectrum of the real signal.
@@ III-B1 Determination of the Number of Signal Components
Assuming the sampling step size of the signal $x(t)$ is $\Delta t$, $x(t)$ can be discretized to $x(j\Delta t)=x_{j}$. For a data window with a length of $T_{\rm w}$, which contains $N=2n+1$ samples. The Hankel matrix $\boldsymbol{X}$ composed of signals within this data window can be represented as [5]:
@@
where every two consecutive singular values correspond to a signal component and the singular values corresponding to signal components are much greater than those corresponding to noises. According to this, the number of signal components $m$ can be determined.
III-B2 Calculation of the Frequency of Signal Component
The Hankel matrix $\boldsymbol{X}$ can also be decomposed as [6]:
@@ III-B2 Calculation of the Frequency of Signal Component
where the $i$th element in $\boldsymbol{Z}_{\mathrm{L}}$ and $\boldsymbol{Z}_{\mathrm{R}}$ is $z_{i}=\mathrm{e}^{\mathrm{j}2\uppi\frac{f_{i}}{f_{\rm sp}}}$, $z_{i}^{n}=\mathrm{e}^{\mathrm{j}2\uppi n\frac{f_{i}}{f_{\rm sp}}}$, $f_{\rm sp}$ is the sampling frequency, $\boldsymbol{P}$ is a matrix composed of the frequency spectrum of each signal component (referred to as phasor). It is worth noting that $z_{i}$ can be solved based on the matrix pencil techniques [6]. Then, the frequency of each signal is calculated:
@@ III-B3 Calculation of Amplitude and Phase of Each Signal Component
After obtaining $f_{i}$, $\boldsymbol{Z}_{\mathrm{R}}$ and $\boldsymbol{Z}_{\rm L}$ in (10) can be calculated. By using the least squares method to solve (10), the phasor matrix $\boldsymbol{P}$ can also be obtained:
@@
where $\boldsymbol{Z}_{\mathrm{L}}^{\mathrm{H}}$ is the conjugate transposition of $\boldsymbol{Z}_{\mathrm{L}}$. The amplitude ${a}_{i}$ and phase ${\phi}_{i}$ of each signal component can be obtained by:
${a}_{i}=2\left|{p}_{i}\right|,~{}~{}{\phi}_{i}=\angle{p}_{i}$ (13)
@@ III-B4 Analytical Signal Consctruction
When the frequency, amplitude, and phase of each signal component are obtained, the imaginary part signal $T\left(x(t)\right)$ in $\underline{x}(t)$ (denoted as $\hat{x}(t)$) can be accurately calculated:
@@
Based on the original real signal and imaginary part signal, the analytical envelope signal $X(t)$ is obtained:
$X(t)=\underline{x}(t)\mathrm{e}^{-\mathrm{j}\omega_{\rm s}t}=\left(x(t)+\mathrm{j}\hat{x}(t)\right)\mathrm{e}^{-\mathrm{j}\omega_{\rm s}t}$ (15)
It is worth noting that $\underline{x}(t)$ obtained by the ESPRIT is a true analytical signal, which only contains positive-frequency components. Then, the frequencies of the analytical envelope signal $X(t)$ obtained by frequency shift transformation are much lower than that of $x(t)$. On the contrary, the analytical signals obtained by methods of the delay transformation [3] and the second parallel simulation [4] are virtual analytical signals. When there are harmonic components, if the methods in [3] and [4] are used, the analytical envelope signal $X(t)$ in the interface will contain signal components with higher frequencies than the original signal.
@@ IV Case Studies
The effectiveness of the proposed method is validated on a modified IEEE 39-node system with a wind farm, whose topology is shown in Fig. 1. The test system is divided into two subsystems ($\rm S_{1}$ and $\rm S_{2}$). The rated frequency of the ac grid is 50 Hz. According to the parameters of the interconnected transmission lines, the positive and zero traveling time of line 26-28 and 26-29 are calculated: $\tau_{2628}^{\mathrm{p}}=612.13~{}\upmu$s, $\tau_{2628}^{0}=820.99~{}\upmu$s, $\tau_{2629}^{\mathrm{p}}=807.23~{}\upmu$s and $\tau_{2629}^{0}=1.08$ ms.
@@
First, the test system is calculated by the co-simulation with the proposed interface, the delay transformation-based interface [3] and the second parallel simulation-based interface[4], respectively, where $\rm S_{1}$ and $\rm S_{2}$ are calculated by EMT simulation and SFEMT simulation, respectively. The time step-sizes for the two subsystems are 20 $\upmu$s and 500 $\upmu$s, respectively. The current of line 26-25 obtained by different methods is shown in Fig. 2. Due to the integration of wind farm, a subsynchronous oscillation occurs and there are interharmonic components in the current. It can be found that the co-simulation based on the proposed interface is more accurate than the other two interfaces when there are harmonic components. The reason is that the imaginary part of the analytical signal obtained by the methods in [3] and [4] does not satisfy the orthogonal relationship (i.e., (5)) with the original real signal. The analytical signals obtained by these two methods still contain negative frequency components. After frequency shifting, the analytical envelope will contain signal components with higher frequencies. On the contrary, the analytical signal obtained by the proposed method only contains positive-frequency components. After frequency shifting, the frequency of the signal decreases, and the simulation accuracy will thus be improved.
Figure 2: Phase A current of the lines 26-25.
Figure 3: Amplitude spectra of analytical signals obtained by different interfaces. (a) Original signal. (b) Delay transformation. (c) Second parallel simulation. (d) Proposed method.
Next, to further illustrate the advantages of the proposed method, the spectra of the analytical signals obtained from the three different interface methods are compared. For the current of line 26-28 that was obtained by different methods, the amplitude spectra of the analytical signals are shown in Fig. 3. It can be observed that the spectra of the analytical signals obtained with the interfaces of delay transformation [3] or the second parallel simulation [4] contain negative frequency components. In contrast, the spectrum of the analytical signal obtained by the proposed method only contains positive frequency components. After frequency shifting, the obtained analytical envelope signal will become a low-frequency signal, and a larger simulation step size can be used.
@@ V Conclusion
This letter proposes an interface for the co-simulation of EMT and the SFMET models. An analytical signal construction method based on ESPRIT is proposed for the interface, which improves the accuracy of co-simulation. Inspired by this letter, many other techniques in the signal process area may be introduced into the area of power system co-simulation.